OpenStudy (anonymous):

solve this system of equations using substitution put the answer in ordered pair form (x,y). 5x +y = 5 and 3x + 2y = 3

OpenStudy (btaylor):

Solve the first equation (5x+y=5) for y. Now plug that value of y in the second equation. So y=5-5x -->3x+2(5-5x)=3 Now simplify and solve for x. Once you have x, plug that value into either equation, and solve for y.